# Decoupling of neural network calibration measures

### v1

@@ 2.1 Calibration approaches
Since modern neural networks tend to produce overconfident predictions as outputs [11], quantifying the calibration quality (or conversely the calibration error) is essential in perception tasks pertaining to safety-critical domains like autonomous driving or medical imaging [1]. A neural network’s output prediction vector $\mathbf{\hat{p}}(\boldsymbol{y}|\boldsymbol{w},\boldsymbol{x})$, for a multi-class setting with $\boldsymbol{y}\in\mathcal{Y}_{K}$ classes and input instance $\boldsymbol{x}$ as well as trained weights $\boldsymbol{w}$, is interpreted as a confidence vector when the prediction logits from the penultimate layer $\mathcal{L}(\boldsymbol{w},\mathbf{x})$ are passed into a sigmoidal function like the softmax $\sigma$.
@@ 3 Evaluation setup
This work uses a customized UNET architecture [25] with five downsampling stages. The network comprises 7,271,110 trainable parameters and 1,178 non-trainable parameters. Furthermore, the depth of the CNN extends over 48 layers plus the input and output layer. The wideness of the network doubles before every downsample step to conserve information and reaches a maximum of 304 channels at the UNET bottleneck. Additionally, the finesse of the input is given by $604\times 960$ pixels and reaches a minimum in the UNET bottleneck of $19\times 30$ pixels. The weight tensors of the convolutional layers are regularized by kernel orthonormality regularization [14], which encourages the network to learn orthonormal features. Moreover, the loss function is given by the balanced cross-entropy and utilizes a focal loss modulation in order to down-weight the loss assigned to well-classified instances. The cross-entropy is balanced by weighting the classes according to their frequency in the A2D2 dataset, similar to Dreissig et al. [8]. Finally, the UNET has been trained for 400 epochs on the A2D2 segmentation dataset [9] by utilizing two Nvidia RTX A6000 with 48GB. The trained UNET achieves a mIoU of $74.8\%$ on the evaluation set.
@@ 4.2 Class-wise analysis
Class ID Class Reliability Sparsification
@@
7 Nature 0.4 0.4 1.8 0.9
8 Sky 0.4 0.4 1.4 0.9
0 Road 0.5 0.5 1.7 0.1
9 Pedestrian 0.3 0.4 0.9 0.9
12 Small Vehicle 0.2 0.3 1.9 1.0
15 Animal 0.9 1.2 0.7 0.6
Table 1: $T_{\mathrm{cal}}$, which minimizes the respective calibration metric, is listed for the three most represented (top three classes) and the three least represented classes (bottom three classes) in the evaluation set. It is evident that the standard deviation of the class-wise calibration temperature is amplified if a subset of less represented classes is considered.
@@ 4.4 AUSE as an uncertainty estimator
Typically, the cross-entropy is used as a suitable loss function for multiclass semantic segmentation. If an ideal uncertainty measure is available, then the sparsification curve is expected to align with the oracle curve suppressing the AUSE to zero. Hence, we propose that the AUSE, obtained when cross-entropy is employed as the uncertainty measure to sort the predictions, as an estimator for the residual uncertainty. The residual uncertainty is non-reducible if the network architecture is fixed and comprises the aleatoric uncertainty of the underlying data generation process as well as the model uncertainty due to the limitations in the hypothesis space induced by the choice of the neural network architecture. This holds true if the model training converges. It is to be expected that the epistemic uncertainty, given by the sum of the approximation uncertainty and the model uncertainty [13], reduces as model training improves till overfitting starts to kick in. To support this hypothesis, we trained five models with an increasing number of epochs maintaining equivalence w.r.t. the model architecture and the hyperparameters respectively. The most prevalent classes, based on their representation in the chosen evaluation set, are considered for the evaluation of the AUSE utilizing cross-entropy (AUSECE). The AUSE over the number of epochs is plotted in Figure 4 and indicates strictly monotonic decreasing graphs, which supports the hypothesis.
Figure 4: From the evaluation set considered, the classes with an occurrence of $>10^{7}$ are chosen for the analysis. It is evident that the AUSECE reduces to the smallest value for the 400-epoch model for each of these classes.
@@ 5 Conclusion
With our work, we show evidence on the inconsistency between well-established calibration measures, which ultimately prevents the homologation of safety-critical applications based on neural networks. In detail, we studied the dependency of the optimal temperature in temperature scaling for a UNET model on the A2D2 dataset. We found that the optimal temperature retrieved by minimizing the NLL does not match the minimum of the ECE, as it has been previously observed by Nixon et al. [22]. Furthermore, we noticed a similar decoupling between the AUSE and the ECE. This gives rise to a degree of freedom in the calibration of neural networks, which undermines the goal of determining unique prediction intervals. Additionally, we empirically demonstrate that there is no difference in minimizing the ECE or UCE. Both methods lead to the same optimal temperature, which is remarkable keeping in mind that the UCE considers the entire prediction vector and the ECE exclusively relies on the probability mass of the most likely class. In contrast, for the AUSE as a calibration metric, we also noticed a decoupling between the AUSEV, using the point-wise uncertainty quantified by the variation ratio, and the AUSES, utilizing the Shannon entropy based on the entire prediction vector. As a result, current neural network calibration metrics lead to different prediction intervals, which generates an ambiguity that compromises trustworthiness. This conclusion also holds true if the ECE is compared to the CCQS or if the uncertainty-based counterparts, the UCE and the UCQS, are contrasted. For this, we transform the idea of assessing the areal deviation in the predicted confidences versus expected confidences graph in a regression setting to the reliability diagram in a classification problem. In this case, the ambiguity between the neural network calibration measures arises from the skewness of the frequency distribution in the predicted confidences/uncertainties.
In addition, we propose the AUSECE as a measure of the residual uncertainty if the sparsification is achieved by ordering the predictions according to the cross-entropy. If the training of the neural network has converged to the best possible minimum within the loss landscape, the residual uncertainty quantifies the remaining uncertainty originating from the data generation process (aleatoric nature) as well as the model uncertainty due to limitations in the hypothesis space given a fixed network architecture (epistemic nature). As a consequence, the model bias can be isolated by subtracting the conditional entropy, which makes the model uncertainty accessible.

### v2

@@ 2.1 Calibration approaches
Since modern neural networks tend to produce overconfident predictions as outputs [10], quantifying the calibration quality (or conversely the calibration error) is essential in perception tasks pertaining to safety-critical domains like autonomous driving [27] or medical imaging [1]. A neural network’s output prediction vector $\mathbf{\hat{p}}(\boldsymbol{y}|\boldsymbol{w},\boldsymbol{x})$, for a multi-class setting with $\boldsymbol{y}\in\mathcal{Y}_{K}$ classes and input instance $\boldsymbol{x}$ as well as trained weights $\boldsymbol{w}$, is interpreted as a confidence vector when the prediction logits from the penultimate layer $\mathcal{L}(\boldsymbol{w},\mathbf{x})$ are passed into a sigmoidal function like the softmax $\sigma$.
@@ 3 Evaluation setup
This work uses a customized UNET architecture [24] with five downsampling stages. The network comprises 7,271,110 trainable parameters and 1,178 non-trainable parameters, which track the first two statistical moments of the batch distribution during training mode for the batch normalization and are not subject to the loss backpropagation. Furthermore, the depth of the CNN extends over 48 layers plus the input and output layer. The wideness of the network doubles before every downsample step to conserve information and reaches a maximum of 304 channels at the UNET bottleneck. Additionally, the finesse of the input is given by $604\times 960$ pixels and reaches a minimum in the UNET bottleneck of $19\times 30$ pixels. The weight tensors of the convolutional layers are regularized by kernel orthonormality regularization [13], which encourages the network to learn orthonormal features. Moreover, the loss function is given by the balanced cross-entropy and utilizes a focal loss modulation in order to down-weight the loss assigned to well-classified instances. The cross-entropy is balanced by weighting the classes according to their frequency in the A2D2 dataset, similar to Dreissig et al. [7]. Finally, the UNET has been trained for 400 epochs on the A2D2 segmentation dataset [8] by utilizing two Nvidia RTX A6000 with 48GB. The trained UNET achieves a mIoU of $74.8\%$ on the evaluation set.
@@ 4.2 Class-wise analysis
Class Reliability Sparsification
@@
Nature 0.4 $\pm$0.05 0.4 $\pm$0.05 1.8 $\pm$0.17 0.9 $\pm$0.38
Sky 0.4 0.4 1.4 0.9
Road 0.5 0.5 1.7 0.1
Pedestrian 0.3 $\pm$0.31 0.4 $\pm$0.40 0.9 $\pm$0.52 0.9 $\pm$0.17
Small Vehicle 0.2 0.3 1.9 1.0
Animal 0.9 1.2 0.7 0.6
Table 1: $T_{\mathrm{cal}}$, which minimizes the respective calibration metric is listed for the three most and least represented classes in the evaluation set (top three and bottom three respectively). It is noticeable that the standard deviation of the class-wise $T_{\mathrm{cal}}$ is amplified if a subset of less represented classes is considered.
@@ 4.4 AUSE as an uncertainty estimator
Typically, the cross-entropy is used as a suitable loss function for multi-class semantic segmentation. If an ideal uncertainty measure is available, then the sparsification curve is expected to align with the oracle curve suppressing the AUSE to zero. Hence, we propose the AUSE, obtained when cross-entropy is employed as the uncertainty measure to sort the predictions, as an estimator for the residual uncertainty. The residual uncertainty is non-reducible if the network architecture is fixed and comprises the aleatoric uncertainty of the underlying data generation process as well as the model uncertainty due to the limitations in the hypothesis space induced by the choice of the neural network architecture. This holds true if the model training converges. It is to be expected that the epistemic uncertainty, given by the sum of the approximation uncertainty and the model uncertainty [12], reduces as model training improves until overfitting begins to occur. To support this hypothesis, we trained five models with an increasing number of epochs maintaining equivalence w.r.t. the model architecture and the hyperparameters respectively. The most prevalent classes, based on their representation in the chosen evaluation set, are considered for the evaluation of the AUSE utilizing cross-entropy (AUSECE). The AUSE over the number of epochs is plotted in Figure 4 and indicates strictly monotonic decreasing graphs, which supports the hypothesis.
Figure 4: Only classes with an occurrence of $>10^{7}$ are considered. It is evident that the class-wise AUSECE converges to the lowest value for the 400-epoch model.
@@ 5 Conclusion
With our work, we show evidence on the inconsistency between well-established calibration measures, which ultimately prevents the homologation of safety-critical applications based on neural networks. In detail, we studied the dependency of the optimal temperature in temperature scaling for a UNET model on the A2D2 dataset. We found that the optimal temperature retrieved by minimizing the NLL does not match the minimum of the ECE, as observed previously by Nixon et al. [21]. Furthermore, we noticed a similar decoupling between the AUSE and the ECE, indicating a degree of freedom in calibrating neural networks. This undermines the goal of determining unique prediction intervals. We also empirically demonstrate that there is no difference in minimizing the ECE or UCE. Both methods lead to the same optimal temperature, which is remarkable considering that the UCE uses the entire prediction vector and the ECE relies exclusively on the probability mass allocated to the most likely class. In contrast, for the AUSE as a calibration metric, we also noticed a decoupling between the AUSEV, using the point-wise uncertainty quantified by the variation ratio, and the AUSES, utilizing the Shannon entropy based on the entire prediction vector. As a result, current calibration metrics lead to different prediction intervals, generating an ambiguity that compromises trustworthiness. This conclusion also holds true if the ECE is compared to the CCQS or if their uncertainty-based counterparts, the UCE and UCQS are contrasted. For this, we transform the idea of assessing the areal deviation in the predicted confidences versus expected confidences graph in a regression setting to the reliability diagram in a classification problem. The ambiguity between these calibration measures arises from the skewness of the frequency distribution in the predicted confidences/uncertainties.
In addition, we propose the AUSECE as a measure of the residual uncertainty if predictions are sparsified according to the cross-entropy. If model-training has converged to the best possible minimum within the loss landscape, the residual uncertainty quantifies the remaining uncertainty originating from the data generation process (aleatoric nature) as well as the model uncertainty due to limitations in the hypothesis space given a fixed neural network architecture (epistemic nature). Consequently, the model bias can be isolated by subtracting the conditional entropy, which makes the model uncertainty accessible.